# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v1

@@ III-C Proof of Theorem 3 (Converse for SUPERSET)
(Adapted from [17, Propositions 3.4 and 3.5])  Let $M_{1}$ denote the number of masked defective items, where there are $k=\Theta(n^{\theta})$ defective items present within a population of size $n$. Let $\xi>0$ be arbitrarily small, and fix $\varepsilon>0$ and $\theta^{\prime}$ sufficiently close to $1$ such that $2(1-\theta^{\prime})<\xi<\varepsilon\theta^{\prime}$. Then if $\theta\in\big{(}\frac{\log 2}{1+\log 2},\frac{1}{2}\big{)}$ and $T\leq\big{(}\frac{1}{\log 2}k\log_{2}k\big{)}(1-\varepsilon)$, it holds that
@@
Defining the event =$\mathscr{B}=\{k\leq\lvert S^{\prime}\rvert\leq k+2\sqrt{k}\log n\}$, it follows due to the multiplicative Chernoff bound (see Appendix -A) that $\mathbb{P}(\mathscr{B})=1-o(1)$. Moreover, conditioned on $\lvert S^{\prime}\rvert\geq k$ (which is implied by $\mathscr{B}$), the resulting distribution of $S$ is indeed the combinatorial prior due to the symmetry of the construction. Thus, an event has $o(1)$ probability under the above distribution if and only if it has $o(1)$ probability under the combinatorial prior. Additionally, letting $M_{1}^{\prime}$ is the number of masked items after the first step of the construction [^5] Specifically, we use this terminology to mean the number of defectives that would be masked if $S^{\prime}$ were the defective set. and $M_{1}$ denote the number after the second step, it holds that $M_{1}^{\prime}\geq M_{1}$. This is because converting defective items to non-defective in step 2 can only decrease the number of masked items, since items that were masked by such “converted” items may no longer be masked.
@@
We proceed by showing that ${\mathcal{A}}_{1}=o(1)$, which in turn implies ${\mathcal{A}}_{2}=1-o(1)$. Under the i.i.d. prior, it is known that the posterior distribution of masked items matches the prior [29, Lemma 3.1]. Thus, since the first step of the generation of $S$ is i.i.d. with parameter $q^{\prime}$, it holds that $(M_{1}^{\prime}\mid M^{\prime}=m^{\prime})\sim\mathrm{Bin}(m^{\prime},q^{\prime})$. Substituting the choice of $q^{\prime}$, the mean is given by $\mu^{\prime}_{1}\coloneqq\mathbb{E}[M_{1}^{\prime}\mid M^{\prime}=m^{\prime}]=m^{\prime}\frac{k+\sqrt{k}\log n}{n}\leq k^{\frac{\gamma}{2}}\big{(}1+o(1)\big{)}$ when $m^{\prime}\leq\frac{n}{k}k^{\frac{\gamma}{2}}$. Thus, writing $\frac{1}{4}k^{\gamma}$ in the form $(1+\delta_{4})\mu^{\prime}_{1}$ for some $\delta_{4}>0$ implies that $\delta_{4}=\Omega(k^{\frac{\gamma}{2}})$. We thus have from the multiplicative Chernoff bound (Appendix -A) that
@@
  • (63) follows since for an item to no longer be masked after the second step, there needs to be at least one test such that all the defective items masking it are converted to non-defective.
  • (64) follows from the union bound.
  • (65) follows by bounding the probability that all $j\in{\mathcal{M}}_{t}^{\prime}(i)$ are converted to non-defective by the probability of this happening for a single $j$ (e.g., the one with the lowest index). We can then bound the probability of this single item being converted by $\frac{2\sqrt{k}\log n}{k}$, since on $\mathscr{B}$ at most $2\sqrt{k}\log n$ such items get chosen uniformly at random from a set of size at least $k$ to be made non-defective.
@@
Letting $\Delta=M^{\prime}-M$ denote the number of items that lose their masked status after the second step, it follows that $\mathbb{E}\Delta\leq\frac{2\log^{4}n}{\sqrt{k}}M^{\prime}$. Markov’s inequality then gives $\mathbb{P}_{2}(\Delta\geq\frac{\log^{5}n}{\sqrt{k}}M^{\prime})=o(1)$, which implies $M\geq M^{\prime}(1-o(1))$ with probability $1-o(1)$. Combining this with the fact that $M^{\prime}\geq\frac{n}{k}k^{\frac{\gamma}{2}}$ with probability $1-o(1)$, we can thus infer there are at least $\frac{n}{k}k^{\frac{\gamma}{2}}(1-o(1))-k\geq\frac{n}{2k}k^{\frac{\gamma}{2}}=\omega(\frac{n}{k})$ masked non-defective items with probability $1-o(1)$. Since $\theta<\frac{1}{2}$, this quantity is $\omega(k)$.
Dropping the bounded tests-per-item assumption. We now argue that the same finding is true for any test design. Consider an arbitrary test design ${\mathsf{X}}$, and partition the items $[n]$ into $({\mathcal{N}}_{\mathrm{low}},{\mathcal{N}}_{\mathrm{high}})$, where ${\mathcal{N}}_{\mathrm{low}}$ is the set of items that appear in at most $\log^{3}n$ tests, and ${\mathcal{N}}_{\mathrm{high}}$ contains the remaining items. We may assume without loss of generality that no test in ${\mathsf{X}}$ contains more than $n^{1-\theta}\log n$ items, since [17, Lemma 3.7] shows that with probability $1-o(1)$, all tests containing more items are positive (and thus their outcomes could have been guessed reliably without performing them). Under this assumption, a simple degree-counting argument in [17, Lemma 3.8] shows that $\lvert{\mathcal{N}}_{\mathrm{low}}\rvert=n-o(n)$. Additionally, while Lemma 10 is stated for the number of masked defectives, the proof in [17] actually shows this result for masked defectives in ${\mathcal{N}}_{\mathrm{low}}$ (again see [17, Lemma 3.8] and its subsequent steps). In other words, in Lemma 10 we may replace $M_{1}$ by $M_{1}^{\rm low}$, defined as the number of masked defectives in ${\mathcal{N}}_{\mathrm{low}}$. We may then apply steps (52)-(66) with all such quantities (i.e., $M_{1}$, $M^{\prime}_{1}$, $M^{\prime}$, $M$, and the associated sets $\mathcal{M},\mathcal{M}^{\prime}$) replaced by their counterparts that only count items in ${\mathcal{N}}_{\mathrm{low}}$; apart from this modification, the analysis remains identical. It follows that there are at least $\frac{n}{2k}k^{\frac{\gamma}{2}}$ masked non-defectives in ${\mathcal{N}}_{\mathrm{low}}$ with probability $1-o(1)$, which trivially implies the same lower bound on the total number of masked non-defectives.
@@
(Adapted from [29, Eqn. 3.12]) For any $\varepsilon\in(0,1)$, when $k=\Theta(n^{\theta})$ for some $\theta\in\big{[}\frac{1}{2},1\big{)}$ and $T\leq\big{(}\frac{1}{\log 2}k\log_{2}\frac{n}{k}\big{)}(1-\varepsilon)$, the number of masked defectives $M_{1}$ satisfies
@@
  • We can then deduce (similar to (62)) that $M^{\prime}\geq k^{1+\frac{\xi}{2}}$ with probability $1-o(1)$.
@@
  • Finally, we drop the assumption that each item appears in at most $n^{\xi}$ tests; this is done by noting that Lemma 11 still holds when “number of masked items” is replaced by “number of masked items appearing in at most $n^{\xi}$ tests” (see [29, Procedure 3.1, Step 1]) and using the same reasoning as Case 1.
@@
[17, Fact 3.1] For a given test matrix ${\mathsf{X}}$ and outcome vector $\mathbf{y}$, let ${\mathcal{V}}({\mathsf{X}},\mathbf{y})\subseteq{\mathcal{S}}$ be the collection of satisfying sets, i.e. the collection of $s\in{\mathcal{S}}$ that would give rise to the test outcome $\mathbf{y}$ when using the test matrix ${\mathsf{X}}$. Then, under the combinatorial prior, it holds that $(S\mid{\mathsf{X}},\mathbf{y})\sim\mathrm{Unif}({\mathcal{V}}({\mathsf{X}},\mathbf{y}))$.
@@
Denote the plausible sets by $S_{1},\dotsc,S_{N}$ (the dependence on $({\mathsf{X}},\mathbf{y})$ is left implicit), and note that by definition $|S_{i}|=k$. We first argue that for every plausible set, there are many implausible sets obtained by swapping out a single element. Specifically, swapping out one masked defective item (in $S_{i}$) with one masked non-defective item in $[n]\setminus\hat{S}$ produces an implausible set, because being plausible requires being a subset of $\hat{S}$. Note that $S_{i}\subseteq\hat{S}$, so combining $|\hat{S}|=(1+\eta^{+})k$ and $|S_{i}|=k$ gives that there are at most $\eta^{-}k$ masked non-defectives in $\hat{S}$, and hence at least $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k$ masked non-defectives in $[n]\setminus\hat{S}$ (see the definition of $\widetilde{{\mathcal{T}}}$). Thus, there are at least $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k$ ways of performing the swap described above. For convenience, we apply $\eta^{+}=k^{o(1)}$ to simplify $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k\geq\frac{1}{4}k^{1+\frac{\xi}{2}}$ for sufficiently large $k$, and work with this simpler lower bound.
@@
We claim that for each $S_{i}$ chosen, the number of $S_{j}$ ($j\neq i$) removed is at most $\eta^{+}k$, where we recall that $\eta^{+}=k^{o(1)}$. This is because $S_{i}$ and $S_{j}$ necessarily share $k-1$ defectives in common, and there is only one choice for the item in $S_{i}\setminus S_{j}$ (it must be $S_{i}$’s designated defective) and at most $\eta^{+}k$ choices for the item in $S_{j}\setminus S_{i}$ (it must be an element of $\hat{S}\setminus S_{i}$, where $|\hat{S}|=(1+\eta^{+})k$ and $|S_{i}|=k$).
@@
When $\theta\in\big{(}\frac{\log 2}{1+\log 2},\frac{1}{2}\big{)}$, we have established the stronger claim that there are $\omega(\frac{n}{k})$ masked non-defective items with high probability (rather than only $\omega(k)$). We claim that this implies the same result when $\eta^{+}=\Omega(k^{\lambda})$ for $\lambda\in(0,\frac{1}{\theta}-2)$ (rather than only $\eta^{+}=k^{o(1)}$). This is because under such a choice, we have (via $k=\Theta(n^{\theta})$) that the number of possible swaps with items not in $\hat{S}$ is at least $\frac{n}{2k}k^{\frac{\gamma}{2}}-\eta^{+}k=\Theta(\frac{n}{k}k^{\frac{\gamma}{2}})$, and that $\frac{(n/k)k^{\frac{\gamma}{2}}}{\eta^{+}k}\to\infty$, which implies $\frac{\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert}{\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert}=o(1)$ by the same reasoning as that above. ∎
@@ IV Conclusion
Here we provide statements of some standard concentration inequalities (e.g., see [31, Ch. 2]) that we use throughout the paper. Letting $X_{1},X_{2},\dots,X_{n}$ be a sequence of i.i.d. Bernoulli($\mu$) random variables, we have the following:
@@
  • (Weakened Chernoff Bound for Binomial RVs) For any $\delta>0$, we have
@@
Here we outline the proof that algorithm in the algorithm in [15] still succeeds for two-sided approximate recovery when $\beta=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ and $\lambda\in\big{(}0,\frac{1}{2}\big{)}$, as opposed to only constant $\beta\in(0,1)$. We note that a variant of this idea (with $\beta=\Theta(k^{\gamma})$ for $\gamma\in(0,1)$) has been analyzed in [28], but the details differ slightly since they consider the noisy case.
Recall the condition in (10) defining the decoder, and let $TI(\tau)$ be the mean of $\imath^{T}\left({\mathsf{X}}_{s_{\mathrm{dif}}};\mathbf{y}\mid{\mathsf{X}}_{s_{\mathrm{eq}}}\right)$. (Here $I(\tau)$ is a conditional mutual information, but the precise definitions of $\imath^{T}$ and $I(\tau)$ are not needed for our purposes.) The following condition on the number of tests is stated in [15, Eqn. 3.37] for parameters $\delta_{1}$ and $\{\delta_{2,\tau}\}_{\tau=\beta k+1}^{k}$:
@@
Letting ${\mathcal{E}}$ denote the failure event for two-sided approximate recovery, [15, Eqn. 3.38] states that if (80) holds for all $\tau=\beta k+1,\dots,k$, then the error probability is upper bounded as follows:
@@
where $\psi_{\tau}:\mathbb{Z}\times\mathbb{R}\to\mathbb{R}$ comes from a concentration inequality given in [15, Eqn. 3.45], and is defined by
@@
We can directly use part of the analysis in [15, Appendix B], showing that (i) (80) is satisfied for all $\tau=\beta k+1,\dots,k$, and (ii) the second sum in (83) tends to 00 for arbitrarily small $\delta_{2}^{(2)}>0$, as long as $T\geq\big{(}k\log_{2}\frac{n}{k}\big{)}(1+\varepsilon)$. Hence, to show that $\mathbb{P}({\mathcal{E}})\to 0$, it suffices to show that the first sum in (83) also decays to zero. Note that if $\beta k+1>\frac{k}{\log k}$ then the sum is empty and is trivially zero, so henceforth we assume this is not the case. We proceed to bound this first sum as follows via (82):
@@
where (88) follows by taking $\delta_{2}^{(1)}$ to be arbitrarily close to $1$ and introducing $\varepsilon^{\prime}>0$ that can be made arbitrarily small (as a function of $\xi$ and $\delta_{2}^{(1)}$). In order for ${\mathcal{A}}\to 0$, we require that the exponent in (88) approaches $\infty$- ∞. For this to happen, it suffices that the following holds for all $\tau\in\big{\{}\beta k+1,\dotsc,\frac{k}{\log k}\big{\}}$:
@@
By the assumption $\lambda\in(0,\frac{1}{2})$, this condition on $T$ is satisfied when $T=(k\log_{2}\frac{n}{k})(1+\varepsilon)$ for arbitrarily small $\varepsilon>0$. Having established that $\mathbb{P}({\mathcal{E}})\to 0$ under this condition, the proof is complete. ∎
@@
It is shown in [17, Proof of Corollary 3.13] that if $T\leq\big{(}\frac{1}{\log 2}k\log_{2}k\big{)}(1-\varepsilon)$, and if there are $k=\tilde{n}^{\theta^{\prime}}$ defective items in a population of size $\tilde{n}$ for some $\theta^{\prime}$ sufficiently close to one such that $2(1-\theta^{\prime})<\xi<\varepsilon\theta^{\prime}$ for arbitrarily small $\xi>0$, the number of masked defective items $\widetilde{M}_{1}$ is stochastically dominated by $\mathrm{Bin}(\tilde{n}^{1-\xi},\frac{1}{3}\tilde{n}^{\varepsilon\theta^{\prime}-1})$. We then have
@@
Thus, with this parametrization, there are at least $\widetilde{M}_{1}\geq\frac{1}{4}k^{\varepsilon-\frac{\xi}{\theta^{\prime}}}$ masked defective items, with probability $1-o(1)$.
@@
We begin by using [29, Eqn. 3.12] and its preceding paragraph, which state that when using an i.i.d. prior with $\tilde{q}=\frac{k}{\tilde{n}}$ and $\tilde{n}$ items, where $k=\tilde{n}^{\theta^{\prime}}$ for $\theta^{\prime}=1-o(1)$, the number of masked defectives is stochastically dominated by $\mathrm{Bin}(\tilde{n}^{1-3\xi},e^{-\frac{(1+\xi)c_{\tilde{q}}}{\tilde{n}\tilde{q}}T})$ for some constant $c_{\tilde{q}}>0$ that depends on $\tilde{q}$. Moreover, when $\tilde{q}=o(1)$ (which is the case we consider), this constant is given explicitly by $c_{\tilde{q}}=(\log^{2}2)\big{(}1+o(1)\big{)}$. Hence, replacing $\tilde{n}\tilde{q}$ by $k$ in the denominator of the exponent and using standard binomial concentration bounds (Appendix -A), the number of masked items is at least
@@
with probability $1-o(1)$. They then show the same holds for the combinatorial prior in [29, Section 3.2] by viewing the combinatorial prior as being generated in the following two-step procedure:
@@
    Add $\max\{k-\lvert S^{\prime}\rvert,0\}$ items from $[\tilde{n}]\setminus S^{\prime}$ to $S^{\prime}$ uniformly at random to form $S$.

### v3

@@ III-C Proof of Theorem 3 (Converse for SUPERSET)
(Adapted from [16, Propositions 3.4 and 3.5]) Let $M_{1}$ denote the number of masked defective items, where there are $k=\Theta(n^{\theta})$ defective items present within a population of size $n$. Let $\xi>0$ be arbitrarily small, and fix $\varepsilon>0$ and $\theta^{\prime}$ sufficiently close to $1$ such that $2(1-\theta^{\prime})<\xi<\varepsilon\theta^{\prime}$. Then if $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$ and $T\leq\big(\frac{1}{\log 2}k\log_{2}k\big)(1-\varepsilon)$, it holds that
@@
Defining the event $\mathscr{B}=\{k\leq\lvert S^{\prime}\rvert\leq k+2\sqrt{k}\log n\}$, it follows due to the multiplicative Chernoff bound (see Appendix -A) that $\mathbb{P}(\mathscr{B})=1-o(1)$. Moreover, conditioned on $\lvert S^{\prime}\rvert\geq k$ (which is implied by $\mathscr{B}$), the resulting distribution of $S$ is indeed the combinatorial prior due to the symmetry of the construction. Thus, an event has $o(1)$ probability under the above distribution if and only if it has $o(1)$ probability under the combinatorial prior. Additionally, letting $M_{1}^{\prime}$ is the number of masked items after the first step of the construction [^5] Specifically, we use this terminology to mean the number of defectives that would be masked if $S^{\prime}$ were the defective set. and $M_{1}$ denote the number after the second step, it holds that $M_{1}^{\prime}\geq M_{1}$. This is because converting defective items to non-defective in step 2 can only decrease the number of masked items, since items that were masked by such “converted” items may no longer be masked.
@@
We proceed by showing that ${\mathcal{A}}_{1}=o(1)$, which in turn implies ${\mathcal{A}}_{2}=1-o(1)$. Under the i.i.d. prior, it is known that the posterior distribution of masked items matches the prior [30, Lemma 3.1]. Thus, since the first step of the generation of $S$ is i.i.d. with parameter $q^{\prime}$, it holds that $(M_{1}^{\prime}\mid M^{\prime}=m^{\prime})\sim\mathrm{Bin}(m^{\prime},q^{\prime})$. Substituting the choice of $q^{\prime}$, the mean is given by $\mu^{\prime}_{1}\coloneqq\mathbb{E}[M_{1}^{\prime}\mid M^{\prime}=m^{\prime}]=m^{\prime}\frac{k+\sqrt{k}\log n}{n}\leq k^{\frac{\gamma}{2}}\big(1+o(1)\big)$ when $m^{\prime}\leq\frac{n}{k}k^{\frac{\gamma}{2}}$. Thus, writing $\frac{1}{4}k^{\gamma}$ in the form $(1+\delta_{4})\mu^{\prime}_{1}$ for some $\delta_{4}>0$ implies that $\delta_{4}=\Omega(k^{\frac{\gamma}{2}})$. We thus have from the multiplicative Chernoff bound (Appendix -A) that
@@
  • (63) follows since for an item to no longer be masked after the second step, there needs to be at least one test such that all the defective items masking it are converted to non-defective;
  • (64) follows from the union bound;
  • (65) follows by bounding the probability that all $j\in{\mathcal{M}}_{t}^{\prime}(i)$ are converted to non-defective by the probability of this happening for a single $j$ (e.g., the one with the lowest index). We can then bound the probability of this single item being converted by $\frac{2\sqrt{k}\log n}{k}$, since on $\mathscr{B}$ at most $2\sqrt{k}\log n$ such items get chosen uniformly at random from a set of size at least $k$ to be made non-defective;
@@
Letting $\Delta=M^{\prime}-M$ denote the number of items that lose their masked status after the second step, it follows that $\mathbb{E}\Delta\leq\frac{2\log^{4}n}{\sqrt{k}}M^{\prime}$. Markov’s inequality then gives $\mathbb{P}_{2}(\Delta\geq\frac{\log^{5}n}{\sqrt{k}}M^{\prime})=o(1)$, which implies $M\geq M^{\prime}(1-o(1))$ with probability 1o(1)1-o(1). Combining this with the fact that $M^{\prime}\geq\frac{n}{k}k^{\frac{\gamma}{2}}$ with probability 1o(1)1-o(1), we can thus infer there are at least $\frac{n}{k}k^{\frac{\gamma}{2}}(1-o(1))-k\geq\frac{n}{2k}k^{\frac{\gamma}{2}}=\omega(\frac{n}{k})$ masked non-defective items with probability 1o(1)1-o(1). Since $\theta<\frac{1}{2}$, this quantity is $\omega(k)$.
Dropping the bounded tests-per-item assumption. We now argue that the same finding is true for any test design. Consider an arbitrary test design $\mathsf{X}$}, and partition the items $[n]$ into $({\mathcal{N}}_{\mathrm{low}},{\mathcal{N}}_{\mathrm{high}})$, where ${\mathcal{N}}_{\mathrm{low}}$ is the set of items that appear in at most $\log^{3}n$ tests, and ${\mathcal{N}}_{\mathrm{high}}$ contains the remaining items. We may assume without loss of generality that no test in $\mathsf{X}$} contains more than $n^{1-\theta}\log n$ items, since [16, Lemma 3.7] shows that with probability 1o(1)1-o(1), all tests containing more items are positive (and thus their outcomes could have been guessed reliably without performing them). Under this assumption, a simple degree-counting argument in [16, Lemma 3.8] shows that $\lvert{\mathcal{N}}_{\mathrm{low}}\rvert=n-o(n)$. Additionally, while Lemma 10 is stated for the number of masked defectives, the proof in [16] actually shows this result for masked defectives in ${\mathcal{N}}_{\mathrm{low}}$ (again see [16, Lemma 3.8] and its subsequent steps). In other words, in Lemma 10 we may replace $M_{1}$ by $M_{1}^{\rm low}$, defined as the number of masked defectives in ${\mathcal{N}}_{\mathrm{low}}$. We may then apply steps (52)-(66) with all such quantities (i.e., $M_{1}$, $^{\prime}_{1}, MM^{\prime}$, $M$, and the associated sets $\mathcal{M},\mathcal{M}^{\prime}$) replaced by their counterparts that only count items in ${\mathcal{N}}_{\mathrm{low}}$; apart from this modification, the analysis remains identical. It follows that there are at least $\frac{n}{2k}k^{\frac{\gamma}{2}}$ masked non-defectives in ${\mathcal{N}}_{\mathrm{low}}$ with probability 1o(1)1-o(1), which trivially implies the same lower bound on the total number of masked non-defectives.
@@
(Adapted from [30, Eqn. 3.12]) For any $\varepsilon\in(0,1)$, when $k=\Theta(n^{\theta})$ for some $\theta\in\big[\frac{1}{2},1\big)$ and $T\leq\big(\frac{1}{\log 2}k\log_{2}\frac{n}{k}\big)(1-\varepsilon)$, the number of masked defectives $M_{1}$ satisfies
@@
  • We can then deduce (similar to (62)) that $M^{\prime}\geq k^{1+\frac{\xi}{2}}$ with probability 1o(1)1-o(1).
@@
  • Finally, we drop the assumption that each item appears in at most $n^{\xi}$ tests; this is done by noting that Lemma 11 still holds when “number of masked items” is replaced by “number of masked items appearing in at most $n^{\xi}$ tests” (see [30, Procedure 3.1, Step 1]) and using the same reasoning as Case 1.
@@
[16, Fact 3.1] For a given test matrix $\mathsf{X}$} and outcome vector $\mathbf{y}$, let ${\mathcal{V}}({\mathsf{X}},\mathbf{y})\subseteq{\mathcal{S}}$ be the collection of satisfying sets, i.e. the collection of $s\in{\mathcal{S}}$ that would give rise to the test outcome $\mathbf{y}$ when using the test matrix $\mathsf{X}$}. Then, under the combinatorial prior, it holds that $(S\mid{\mathsf{X}},\mathbf{y})\sim\mathrm{Unif}({\mathcal{V}}({\mathsf{X}},\mathbf{y}))$.
@@
Denote the plausible sets by $S_{1},\dotsc,S_{N}$ (the dependence on $({\mathsf{X}},\mathbf{y})$ is left implicit), and note that by definition $|S_{i}|=k$. We first argue that for every plausible set, there are many implausible sets obtained by swapping out a single element. Specifically, swapping out one masked defective item (in $S_{i}$) with one masked non-defective item in $n]\setminus\hat{S}$ produces an implausible set, because being plausible requires being a subset of $\hat{S}$. Note that $S_{i}\subseteq\hat{S}$, so combining $|\hat{S}|=(1+\eta^{+})k$ and $|S_{i}|=k$ gives that there are at most $\eta^{-}k$ masked non-defectives in $\hat{S}$, and hence at least $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k$ masked non-defectives in $n]\setminus\hat{S}$ (see the definition of $\widetilde{{\mathcal{T}}}$). Thus, there are at least $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k$ ways of performing the swap described above. For convenience, we apply $\eta^{+}=k^{o(1)}$ to simplify $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k\geq\frac{1}{4}k^{1+\frac{\xi}{2}}$ for sufficiently large $k$, and work with this simpler lower bound.
@@
We claim that for each $S_{i}$ chosen, the number of $S_{j}$ ($j\neq i$) removed is at most $\eta^{+}k$, where we recall that $\eta^{+}=k^{o(1)}$. This is because $S_{i}$ and $S_{j}$ necessarily share k1k-1 defectives in common, and there is only one choice for the item in $S_{i}\setminus S_{j}$ (it must be $S_{i}$’s designated defective) and at most $\eta^{+}k$ choices for the item in $S_{j}\setminus S_{i}$ (it must be an element of $\hat{S}\setminus S_{i}$, where $|\hat{S}|=(1+\eta^{+})k$ and $|S_{i}|=k$).
@@
When $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$, we have established the stronger claim that there are $\omega(\frac{n}{k})$ masked non-defective items with high probability (rather than only $\omega(k)$). We claim that this implies the same result when $\eta^{+}=\Omega(k^{\lambda})$ for $\lambda\in[0,\frac{1}{\theta}-2)$ (rather than only $\eta^{+}=k^{o(1)}$). This is because under such a choice, we have (via $k=\Theta(n^{\theta})$) that the number of possible swaps with items not in $\hat{S}$ is at least $\frac{n}{2k}k^{\frac{\gamma}{2}}-\eta^{+}k=\Theta(\frac{n}{k}k^{\frac{\gamma}{2}})$, and that $\frac{(n/k)k^{\frac{\gamma}{2}}}{\eta^{+}k}\to\infty$, which implies $\frac{\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert}{\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert}=o(1)$ by the same reasoning as that above. ∎
@@ IV Conclusion
Here we provide statements of some standard concentration inequalities (e.g., see [32, Ch. 2]) that we use throughout the paper. Letting $X_{1},X_{2},\dots,X_{n}$ be a sequence of i.i.d. Bernoulli($\mu$) random variables, we have the following:
@@
  • (Weakened Chernoff Bound for Binomial RVs) For any $\delta\in(0,1]$, we have
@@
Here we outline the proof that algorithm in the algorithm in [18] still succeeds for two-sided approximate recovery when $\beta=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ and $\lambda\in\big[0,\frac{1}{2}\big)$, as opposed to only constant $\beta\in(0,1)$. We note that a variant of this idea (with $\beta=\Theta(k^{\gamma})$ for $\gamma\in(0,1)$) has been analyzed in [29], but the details differ slightly since they consider the noisy case.
Recall the condition in (10) defining the decoder, and let $TI(\tau)$ be the mean of $\imath^{T}\left({\mathsf{X}}_{s_{\mathrm{dif}}};\mathbf{y}\mid{\mathsf{X}}_{s_{\mathrm{eq}}}\right)$. (Here $I(\tau)$ is a conditional mutual information, but the precise definitions of $\imath^{T}$ and $I(\tau)$ are not needed for our purposes.) The following condition on the number of tests is stated in [18, Eqn. 3.37] for parameters $\delta_{1}$ and $\{\delta_{2,\tau}\}_{\tau=\beta k+1}^{k}$:
@@
Letting $\mathcal{E}$} denote the failure event for two-sided approximate recovery, [18, Eqn. 3.38] states that if (80) holds for all $\tau=\beta k+1,\dots,k$, then the error probability is upper bounded as follows:
@@
where $\psi_{\tau}:\mathbb{Z}\times\mathbb{R}\to\mathbb{R}$ comes from a concentration inequality given in [18, Eqn. 3.45], and is defined by
@@
We can directly use part of the analysis in [18, Appendix B], showing that (i) (80) is satisfied for all $\tau=\beta k+1,\dots,k$, and (ii) the second sum in (83) tends to 0 for arbitrarily small $\delta_{2}^{(2)}>0$, as long as $T\geq\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$. Hence, to show that $\mathbb{P}({\mathcal{E}})\to 0$, it suffices to show that the first sum in (83) also decays to zero. Note that if $\beta k+1>\frac{k}{\log k}$ then the sum is empty and is trivially zero, so henceforth we assume this is not the case. We proceed to bound this first sum as follows via (82):
@@
where (88) follows by taking $\delta_{2}^{(1)}$ to be arbitrarily close to $1$ and introducing $\varepsilon^{\prime}>0$ that can be made arbitrarily small (as a function of $\xi$ and $\delta_{2}^{(1)}$). In order for ${\mathcal{A}}\to 0$, we require that the exponent in (88) approaches -$\infty$. For this to happen, it suffices that the following holds for all $\tau\in\big\{\beta k+1,\dotsc,\frac{k}{\log k}\big\}$:
@@
By the assumption $\lambda\in[0,\frac{1}{2})$, this condition on $T$ is satisfied when $T=(k\log_{2}\frac{n}{k})(1+\varepsilon)$ for arbitrarily small $\varepsilon>0$. Having established that $\mathbb{P}({\mathcal{E}})\to 0$ under this condition, the proof is complete. ∎
@@
It is shown in [16, Proof of Corollary 3.13] that if $T\leq\big(\frac{1}{\log 2}k\log_{2}k\big)(1-\varepsilon)$, and if there are $k=\tilde{n}^{\theta^{\prime}}$ defective items in a population of size $\tilde{n}$ for some $\theta^{\prime}$ sufficiently close to one such that $2(1-\theta^{\prime})<\xi<\varepsilon\theta^{\prime}$ for arbitrarily small $\xi>0$, the number of masked defective items $\widetilde{M}_{1}$ is stochastically dominated by $\mathrm{Bin}(\tilde{n}^{1-\xi},\frac{1}{3}\tilde{n}^{\varepsilon\theta^{\prime}-1})$. We then have
@@
Thus, with this parametrization, there are at least $\widetilde{M}_{1}\geq\frac{1}{4}k^{\varepsilon-\frac{\xi}{\theta^{\prime}}}$ masked defective items, with probability 1o(1)1-o(1).
@@
We begin by using [30, Eqn. 3.12] and its preceding paragraph, which state that when using an i.i.d. prior with $\tilde{q}=\frac{k}{\tilde{n}}$ and $\tilde{n}$ items, where $k=\tilde{n}^{\theta^{\prime}}$ for $\theta^{\prime}=1-o(1)$, the number of masked defectives is stochastically dominated by $\mathrm{Bin}(\tilde{n}^{1-3\xi},e^{-\frac{(1+\xi)c_{\tilde{q}}}{\tilde{n}\tilde{q}}T})$ for some constant $c_{\tilde{q}}>0$ that depends on $\tilde{q}$. Moreover, when $\tilde{q}=o(1)$ (which is the case we consider), this constant is given explicitly by $c_{\tilde{q}}=(\log^{2}2)\big(1+o(1)\big)$. Hence, replacing $\tilde{n}\tilde{q}$ by $k$ in the denominator of the exponent and using standard binomial concentration bounds (Appendix -A), the number of masked items is at least
@@
with probability 1o(1)1-o(1). They then show the same holds for the combinatorial prior in [30, Section 3.2] by viewing the combinatorial prior as being generated in the following two-step procedure:
@@
    Add $\max\{k-\lvert S^{\prime}\rvert,0\}$ items from $\tilde{n}]\setminus S^{\prime}$ to $S^{\prime}$ uniformly at random to form $S$.